ryHelp & Support Georgia Crag estion 3 t yec arnered arked out of Following is an extract of information from the books of Express Photos 2016 2015 Net Sales Net Credit Sales 15,000 otal Assets urrent Liabilites 80,500 77,000 Gross Profit urrent Assets Net Profit Total Liabilities Accounts Receivable 56, 18, Non leap year Calculate the following ratios Format as noted Current Ratio (to 1 decimal place) Debt Ratio (as a percentage to 1 decimal place-no%) Days Sales in Receivables (rounded to nearest whole number) Inventory Turnover rounded to nearest whole number) Rate if Return on Net Sales (as a percentage to 1 decimal place-no w Select the ratio that would generally be considered better for each of the following
ibraryHelp & Support Georgia Craig kecevapie eap year Calculate the following ratios: Format as noted Current Ratio (to 1 decimal place) Debt Ratio (as a percentage to 1 decimal place-no) Days Sales in Receivables (rounded to nearest whole number) Inventory Turnover (rounded to nearest whole number) Rate if Return on Net Sales (as a percentage to 1 decimal place no Select the ratio that would generally be considered better for each of the following: Current Ratio 1.5 Debt Ratio 6596 Inventory Turnover Gross Profit Accounts Receivable Turnover 20 65% : 5 8
Answer #1

1) Current Ratio = Current Assets / Current Liabilities

= 77000 / 48000 = 1.6

2) Debt Ratio = Total Liabilities / Total Assets

= 56000 / 98000 = 57.1

3) Days Sales in Receivables = Accounts Receivable / Net Credit Sales * 365

= 25000 / 150500 * 365 = 61

4) Inventory Turnover Ratio = Cost of goods sold / Average Inventory

= 86000 / 14000 = 6 times

** Average Inventory = Opening Inventory + Closing Inventory / 2

= 15000 + 13000 / 2 = 14000

5) Return on Net Sales Ratio = Operating Profit / Net Sales

= 55000 / 166500 = 33.03

* Net Profit is assumed as Operating Profit.